# Characterization of fractional Chern insulator quasiparticles in moiré transition metal dichalcogenides

### v1

@@ Abstract
We provide a detailed study of Abelian quasiparticles of valley polarized fractional Chern insulators (FCIs) residing in the top valence band of twisted bilayer MoTe2 (tMoTe2) at hole filling $\nu_{h}=2/3$. We construct a tight-binding model of delocalized quasiparticles to capture the energy dispersion of a single quasiparticle. We then localize quasiparticles by short-range delta impurity potentials. Unlike the fractional quantum Hall (FQH) counterpart in the lowest Landau level (LLL), the density profile around the localized FCI quasiparticle in tMoTe2 depends on the location of the impurity potential and loses the continuous rotation invariance. The FCI quasiparticle localized at moiré lattice center closely follows the anyon Wannier state of the tight-binding model of the mobile quasiparticle. Despite of the difference in density profiles, we find that the excess charge around the impurity potential for the $\nu_{h}=2/3$ FCIs in tMoTe2 is still similar to that of the $\nu=2/3$ FQH state in the LLL if an effective magnetic length on the moiré lattice is chosen as the length unit, which allows a rough estimation of the spatial extent of the FCI quasiparticle. Far away from the impurity potential, this excess charge has the tendency to reach $e/3$, as expected for the Laughlin quasiparticle. The braiding phase of two FCI quasiparticles in tMoTe2 also agrees with the theoretical prediction of fractional statistics. Based on the nearly ideal quantum geometry of the top valence band of tMoTe2, we propose a trial wave function for localized FCI quasiparticles, which reproduces the key feature of the density profile around a quasiparticle. We also discuss the effect of band mixing on FCI quasiparticles in tMoTe2.
@@ I Introduction
The nontrivial topology and enhanced electron-electron interactions in two-dimensional narrow bands with nonzero Chern number have given rise to various novel correlated phases of matters. As a representative example, interacting particles partially occupying a nearly flat topological band can form zero-field analogues of the celebrated fractional quantum Hall (FQH) effect [1, 2] at suitable filling factors, dubbed fractional Chern insulators (FCIs)  [3, 4, 5, 6, 7, 8, 9, 10]. A main trend in the recent development of this field is to gather inspiration from material science to look for lattice platforms that can really produce FCIs. Remarkably, FCIs were theoretically predicted to exist in van-der-Waals heterostructures with moiré patterns [11, 12, 13, 14, 15, 16, 17]. Consistent with these theoretical predictions, experiments have reported observations of FCIs in rhombohedral graphene aligned with hexagonal boron nitride [18, 19, 20, 21, 22, 23] and twisted homobilayer MoTe2 (tMoTe2[24, 25, 26, 27, 28, 29, 30, 31]. Further theoretical studies flourish to understand the exciting experimental results [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50].
@@
In this work, we present a detailed numerical characterization of the Abelian quasiparticles of $\nu_{h}=2/3$ FCIs in tMoTe2. We adopt the continuum model of tMoTe2 and focus on experimentally realistic twist angles near $4^{\circ}$. In the presence of interactions, the many-body calculations are mostly carried out using extensive exact diagonalization (ED) under the assumption of valley polarization and single-band projection. First of all, we find that the hole’s real-space density of the FCI ground state, unlike its FQH counterpart, is strongly modulated within a moiré unit cell, while the total hole numbers in different unit cells remains identical. The density non-uniformness can be interpreted by the confinement of holes in specific regions of the moiré lattice. We then create quasiparticles in the FCI ground state and construct an effective tight-binding model to describe the dispersion of a single mobile quasiparticle. This dispersion, together with the spectrum of two mobile quasiparticles, can be used to estimate the interaction between two quasiparticles, which may predict whether bunches of two quasiparticles occurs. Supposing holes interact via the weakly screened Coulomb potential, we find signature of repulsive interactions between mobile quasiparticles.
[…]
@@ II FQH quasiparticles at $\nu=2/3$ in the LLL
Before studying the $\nu_{h}=2/3$ FCI quasiparticles in tMoTe2, we first characterize quasiparticles of the conventional FQH state at $\nu=2/3$ in the LLL, for the sake of comparing quasiparticle physics in moiré flat bands and in the LLL. Note that we will focus on $\nu=2/3$ quasiparticles rather than quasiholes. We find that ED cannot deal with the latter in finite systems due to its too large spatial extent.
@@ II.1 Model of the LLL
We consider $N$ interacting fermions of charge $e$ on a square torus of length $L$ penetrated by a uniform perpendicular magnetic field $B$. The number of flux quanta, $N_{\phi}$, through the surface of the torus is an integer satisfying $L^{2}=2\pi\ell_{B}^{2}N_{\phi}$, where $\ell_{B}=\sqrt{\hbar/(eB)}$ is the magnetic length. The filling factor is defined as $\nu=N/N_{\phi}$. In a strong magnetic field, we can restrict all fermions in the LLL, such that the kinetic energy is simply a constant. In this case, the effective part in the many-body Hamiltonian is just the (two-body) interaction between fermions, which takes the form of
@@
Here $c_{m}^{\dagger}$ ($c_{m}$) creates (annihilates) a fermion in the $m$th LLL orbital, whose wave function is $\psi_{m}(x,y)=\left(\frac{1}{\sqrt{\pi}L\ell_{B}}\right)^{1/2}\sum_{n=-\infty}^{+\infty}e^{i\frac{2\pi}{L}(m+nN_{\phi})y}e^{-\frac{1}{2\ell_{B}^{2}}\left[x-\frac{2\pi}{L}\ell_{B}^{2}(m+nN_{\phi})\right]^{2}}$ under the Landau gauge. The interaction matrix element is
@@
×$\times V({\bm{q}})e^{-\frac{1}{2}|{\bm{q}}|^{2}\ell_{B}^{2}}e^{i\frac{2\pi s}{N_{\phi}}(m_{1}-m_{3})},$ (2)
@@
where $\delta_{i,j}^{{\rm mod}\ N_{\phi}}$ is the periodic Kronecker delta function with period $N_{\phi}$, ${\bm{q}}=(q_{x},q_{y})=\frac{2\pi}{L}(s,t)$, and $V({\bm{q}})$ is the Fourier transform of the interaction potential. We consider the realistic dual-gate setup and choose the screened Coulomb potential $V({\bm{q}})=\frac{e^{2}}{4\pi\epsilon\epsilon_{0}}\frac{1}{L^{2}}\frac{2\pi\tanh(|{\bm{q}}|d)}{|{\bm{q}}|}$ for the interaction, where $\epsilon$ is the relative dielectric constant of the material and $d$ is the distance from the top (or bottom) gate to the two-dimensional electron gas. We set $d=5\ell_{B}$ (weak screening) throughout our calculations in this section. The results are not sensitive to the precise value of $d$ so long as the screening is not too strong.
@@
The ground state of the gate-screened Coulomb interaction at $\nu=2/3$ is approximately the particle-hole conjugate of the celebrated Laughlin model state, with three-fold topological degeneracy on the torus and Abelian quasiparticle excitations. To generate $N_{\rm qp}$ quasiparticles in the ground state, we add $N_{\rm qp}$ magnetic fluxes and $N_{\rm qp}$ fermions to the system, leading to $N_{\phi}=\frac{1}{2}(3N-N_{\rm qp})$. We consider the cases with sufficiently dilute quasiparticles, such that the charge gap of the finite system remains open. In principle, we should diagonalize the full Hamiltonian $H_{\rm int}+\sum_{j=1}^{N_{\rm qp}}U_{\rm imp}({\bm{w}}_{j})$ to get states with localized quasiparticles, where $U_{\rm imp}({\bm{w}})$ is an attractive impurity potential aiming to pin a quasiparticle at the specific position ${\bm{w}}$. However, localized quasiparticles break the translation invariance on the torus, making this direct diagonalization computationally expensive.
@@
Fortunately, as long as the the charge gap of the system remains finite, the energy spectrum of $H_{\rm int}$ always contains a low-energy subspace, corresponding to $N_{\rm qp}$ delocalized quasiparticles. The dimension of this subspace can be predicted by Haldane’s exclusion principle [83] or the conformal field theory [84]. Assuming that the impurity potentials only mix states within this low-energy subspace without coupling them to higher excited levels, we first diagonalize $H_{\rm int}$ to obtain the states of delocalized quasiparticles, then diagonalize the impurity potentials in the corresponding subspace to localize quasiparticles. The advantage of this two-step method is that the translation invariance is still available in the first step to reduce the many-body Hilbert space dimension. For the $\nu=2/3$ Coulomb ground state, we find an attractive $\delta$ impurity potential $U_{\rm imp}({\bm{w}})=-U_{0}\sum_{i=1}^{N}\delta({\bm{r}}_{i}-{\bm{w}})$, where ${\bm{r}}_{i}$ is the position of the $i$th fermion and $U_{0}>0$, can readily pin a quasiparticle at ${\bm{w}}$. The second quantization form of such an impurity potential projected to the LLL is
@@ II.2 Single quasiparticle
Figure 1: (a) The density distribution for a single-quasiparticle excitation of the $\nu=2/3$ fermionic FQH state on a square torus with $N=25,N_{s}=37$. (b) The excess charge around the impurity potential. The horizontal reference line indicates $Q=e/3$. The dashed line in (b) is obtained using the first-order Haldane’s pseudopotential. Otherwise the interaction is the screened Coulomb potential.
In Fig. 1, we display the results for the largest system size, $N=25,N_{\phi}=37$, that we can reach using ED. There is a pronounced peak at the location of the impurity potential [Fig. 1(a)]. To verify that the density peak corresponds to a localized quasiparticle, we compute the excess charge defined as
@@
[…]
@@
[…]
where $r_{\max}$ is the largest distance from the torus center. A numerical calculation of Eq. (5) for the system size in Fig. 1 shows $R_{\rm qp}\approx 4\ell_{B}$. If we replace the Coulomb interaction with the short-range first-order Haldane’s pseudopotential [86], we get $R_{\rm qp}\approx 2.6\ell_{B}$. The smaller spatial extent of the quasiparticle for the pseudopotential interaction can also be seen in the excess charge [dashed lines in Fig. 1(b)], which almost converges to $e/3$ for the same system size.
@@ II.3 Braiding of two quasiparticles
The total Berry phase can be split into two parts: one is the Aharonov-Bohm (AB) phase caused by moving a single quasiparticle in the uniform magnetic field along the same path without the other quasiparticle enclosed; and the other comes from the fractional statistics between two quasiparticles. The AB phase $p_{i}^{\rm AB}$ should be $(\pi/3)(D/\ell_{B})^{2}$, which we have numerically confirmed. Then we obtain the three fractional statistical phases as $p_{i}^{\rm anyon}=p_{i}-p_{i}^{\rm AB}$. Their averaged value $\phi_{\rm anyon}$ is displayed in Fig. 2 as a function of $D$ for $N=24,N_{\phi}=35$, which is the largest system size we can reach in the presence of two quasiparticles.
When $D$ is large enough, we expect that $\phi_{\rm anyon}$ exponentially converges to the theoretical value $34\pi/$ 3 (equivalently, $-2\pi/3$). However, as shown in Fig. 2, the deviation from this theoretical value still oscillates around $10^{-2}\pi$ even in the largest system size for the farthest separations between two quasiparticles. Therefore, bigger samples are required to further reduce this deviation. The critical value of $D$ at which $\phi_{\rm anyon}$ is close enough to the theoretical value can be used as another estimation of quasiparticle size. We set a threshold $10^{-2}$ for the revival in the oscillations of $|\phi_{\rm anyon}/\pi-4/3|$. Using an exponential fitting for the first two revivals, we predict the critical $D\approx 12\ell_{B}$, which gives the quasiparticle radius as $R_{\rm qp}\approx 6\ell_{B}$. Compared with the Coulomb case, the pseudopotential leads to faster convergence of the fractional statistical phase (Fig. 2), suggesting $R_{\rm qp}\approx 3.5\ell_{B}$.

### v3

# Characterization of fractional Chern insulator quasiparticles in twisted homobilayer MoTe2

@@ Abstract
We provide a detailed study of Abelian quasiparticles of valley polarized fractional Chern insulators (FCIs) residing in the top valence band of twisted bilayer MoTe2 (tMoTe2) at hole filling $\nu_{h}=2/3$. We construct a tight-binding model of delocalized quasiparticles to capture the energy dispersion of a single quasiparticle. We then localize quasiparticles by short-range delta impurity potentials. Unlike the fractional quantum Hall (FQH) counterpart in the lowest Landau level (LLL), the density profile around the localized FCI quasiparticle in tMoTe2 depends on the location of the impurity potential and loses the continuous rotation invariance. The FCI quasiparticle localized at moiré lattice center closely follows the anyon Wannier state of the tight-binding model of the mobile quasiparticle. Despite of the difference in density profiles, we find that the excess charge around the impurity potential for the $\nu_{h}=2/3$ FCIs in tMoTe2 is still similar to that of the $\nu=2/3$ FQH state in the LLL if an effective magnetic length on the moiré lattice is chosen as the length unit, which allows a rough estimation of the spatial extent of the FCI quasiparticle. Far away from the impurity potential, this excess charge has the tendency to reach $e/3$, as expected for the Laughlin quasiparticle. The braiding phase of two FCI quasiparticles in tMoTe2 also agrees with the theoretical prediction of fractional statistics. We characterize the interaction between two FCI quasiparticles and find a crossover from repulsive to attractive interaction as gate-to-sample distances decreases. Based on the nearly ideal quantum geometry of the top valence band of tMoTe2, we propose a trial wave function for localized FCI quasiparticles, which reproduces the key feature of the density profile around a quasiparticle.
@@ I Introduction
The nontrivial topology and enhanced electron-electron interactions in two-dimensional narrow bands with nonzero Chern number have given rise to various novel correlated phases of matters. As a representative example, interacting particles partially occupying a nearly flat topological band can form zero-field analogues of the celebrated fractional quantum Hall (FQH) effect [1, 2] at suitable filling factors, dubbed fractional Chern insulators (FCIs) [3, 4, 5, 6, 7, 8, 9, 10]. A main trend in the recent development of this field is to gather inspiration from material science to look for lattice platforms that can really produce FCIs. Remarkably, FCIs were theoretically predicted to exist in van-der-Waals heterostructures with moiré patterns [11, 12, 13, 14, 15, 16, 17]. Consistent with these theoretical predictions, experiments have reported observations of FCIs in rhombohedral graphene aligned with hexagonal boron nitride [18, 19, 20, 21, 22, 23] and twisted homobilayer MoTe2 (tMoTe2) [24, 25, 26, 27, 28, 29, 30, 31]. Further theoretical studies flourish to understand the exciting experimental results [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54].
@@
In this work, we present a detailed numerical characterization of the Abelian quasiparticles of $\nu_{h}=2/3$ FCIs in tMoTe2. We adopt the continuum model of tMoTe2 and focus on experimentally realistic twist angles near $4^{\circ}$. In the presence of interactions, the many-body calculations are mostly carried out using extensive exact diagonalization (ED) under the assumption of valley polarization and single-band projection. First of all, we find that the hole’s real-space density of the FCI ground state, unlike its FQH counterpart, is strongly modulated within a moiré unit cell, while the total hole numbers in different unit cells remains identical. The density non-uniformness can be interpreted by the confinement of holes in specific regions of the moiré lattice. We then create quasiparticles in the FCI ground state and construct an effective tight-binding model to describe the dispersion of a single mobile quasiparticle. This dispersion, together with the spectrum of two mobile quasiparticles, can be used to estimate the interaction between two quasiparticles, which may predict whether bunches of two quasiparticles occurs. We find the signature of a crossover from repulsive to attractive quasiparticle interaction with the enhancement of screening in the Coulomb potential.
On the other hand, we study the FCI quasiparticles localized by delta impurity potentials. The density profile around a single localized quasiparticle varies with the location of the impurity potential, and is well described by the anyon Wannier states constructed from the tight-binding model of the mobile quasiparticle. The nearly ideal quantum geometry [83, 84, 85, 86] of the top valence band of tMoTe2 allows a variational construction of trial wave functions of FCI anyons. Monte Carlo simulation using these trial wave functions reproduce the key features of the density profile around a localized quasiparticle obtained by ED, thus justifying the validity of trial wave functions. We also estimate the charge of a localized quasiparticle and the braiding statistics of two localized quasiparticles in tMoTe2. The results are consistent with theoretical expectations for the $\nu=2/3$ Laughlin state, i.e., fractional charge of $e/3$ and braiding phase $2\pi/3$. However, the largest system size within the capability of ED is still insufficient for the full development of a quasiparticle. If the effective lattice magnetic length $\ell_{0}\equiv\sqrt{S_{\rm uc}/(2\pi)}$ ($S_{\rm uc}$ is the area of moiré unit cell) [73] and the magnetic length $\ell_{B}$ are used as the length units in tMoTe2 and the lowest Landau level (LLL), respectively, the accumulated excess charge around the impurity potential in tMoTe2 looks similar to that of the $\nu=2/3$ FQH state in the LLL, but shows slower damping with increasing distance. We estimate the quasiparticle radius of the $\nu=2/3$ FQH state in the LLL as $4\sim 6\ell_{B}$ by analyzing the density profile and the braiding phase. This suggests an estimation of the spatial extent of a localized $\nu_{h}=2/3$ FCI quasiparticle in tMoTe2 as $6\ell_{0}$, which is $120\ \mathring$$\mathrm${A}} at twist angle $3.7^{\circ}$. This result could be further increased once band mixing is considered.
@@ II FQH quasiparticles at $\nu=2/3$ in the LLL
Before studying the $\nu_{h}=2/3$ FCI quasiparticles in tMoTe2, we first characterize quasiparticles of the conventional FQH state at $\nu=2/3$ in the LLL, for the sake of comparing quasiparticle physics in moiré flat bands and in the LLL. Note that we will focus on $\nu=2/3$ quasiparticles rather than quasiholes for both FQH and FCI states. We find that localized $\nu=2/3$ quasiholes have too large spatial extent so that ED cannot characterize them well in finite systems (see Appendix E).
@@ II.1 Model of the LLL
We consider $N$ interacting fermions of charge $e$ on a square torus of length $L$ penetrated by a uniform perpendicular magnetic field $B$. The number of flux quanta, $N_{\phi}$, through the surface of the torus is an integer satisfying $L^{2}=2\pi\ell_{B}^{2}N_{\phi}$, where $\ell_{B}=\sqrt{\hbar/(eB)}$ is the magnetic length. As $N_{\phi}$ is also the number of single-particle states (called orbitals) per Landau level, the filling factor is defined as $\nu=N/N_{\phi}$. In a strong magnetic field, we can restrict all fermions in the LLL, such that the kinetic energy is simply a constant. In this case, the effective part in the many-body Hamiltonian is just the (two-body) interaction between fermions, which takes the form of
@@
Here $c_{m}^{\dagger}$ ($c_{m}$) creates (annihilates) a fermion in the $m$th LLL orbital, whose wave function is
$\psi_{m}(x,y)=\left(\frac{1}{\sqrt{\pi}L\ell_{B}}\right)^{1/2}\sum_{n=-\infty}^{+\infty}e^{i\frac{2\pi}{L}(m+nN_{\phi})y}$
$\times e^{-\frac{1}{2\ell_{B}^{2}}\left[x-\frac{2\pi}{L}\ell_{B}^{2}(m+nN_{\phi})\right]^{2}}$ (2)
under the Landau gauge. The interaction matrix element is
@@
$\times V({\bm{q}})e^{-\frac{1}{2}|{\bm{q}}|^{2}\ell_{B}^{2}}e^{i\frac{2\pi s}{N_{\phi}}(m_{1}-m_{3})},$ (3)
@@
where $\delta_{i,j}^{{\rm mod}\ N_{\phi}}$ is the periodic Kronecker delta function with period $N_{\phi}$, ${\bm{q}}=(q_{x},q_{y})=\frac{2\pi}{L}(s,t)$, and $V({\bm{q}})$ is the Fourier transform of the interaction potential. We consider the realistic dual-gate setup and choose the screened Coulomb potential
$V({\bm{q}})=\frac{e^{2}}{4\pi\epsilon\epsilon_{0}}\frac{1}{L^{2}}\frac{2\pi\tanh(|{\bm{q}}|d)}{|{\bm{q}}|}$ (4)
for the interaction, where $\epsilon$ is the relative dielectric constant of the material and $d$ is the distance from the top (or bottom) gate to the two-dimensional electron gas. We set $d=5\ell_{B}$ (weak screening) throughout our calculations in this section. The results are not sensitive to the precise value of $d$ so long as the screening is not too strong.
@@
The ground state of the gate-screened Coulomb interaction at $\nu=2/3$ is approximately the particle-hole conjugate of the celebrated Laughlin model state, with three-fold topological degeneracy on the torus and Abelian quasiparticles. To generate $N_{\rm qp}$ quasiparticles in the ground state, we add $N_{\rm qp}$ magnetic fluxes and $N_{\rm qp}$ fermions to the system, leading to
$N_{\phi}=\frac{1}{2}(3N-N_{\rm qp}).$ (5)
The reason of this choice can be seen from the thin-torus configuration [87] of the $\nu=2/3$ Laughlin state: $110110\cdots 110110$, where $1$ denotes an occupied LLL orbital and 0 denotes an empty LLL orbital. We can create a single $111$ block by adding one occupied orbital (flux), so that the local charge density is increased by $e/3$ compared to the ground state, corresponding to a quasiparticle. The relation between $N_{\phi}$ and $N$ is exactly Eq. (5) if we repeat this procedure for $N_{\rm qp}$ times.
@@
We consider the cases with sufficiently dilute quasiparticles, such that the charge gap of the finite system remains open. In principle, we should diagonalize the full Hamiltonian $H_{\rm int}+\sum_{j=1}^{N_{\rm qp}}U_{\rm imp}({\bm{w}}_{j})$ to get states with localized quasiparticles, where $U_{\rm imp}({\bm{w}})$ is an attractive impurity potential aiming to pin a quasiparticle at the specific position $\bm{w}$}. However, localized quasiparticles break the translation invariance on the torus, making this direct diagonalization computationally expensive. Fortunately, as long as the the charge gap of the system remains finite, the energy spectrum of $H_{\rm int}$ always contains a low-energy subspace, corresponding to $N_{\rm qp}$ delocalized quasiparticles. The dimension of this subspace can be predicted by Haldane’s exclusion principle [88] or the conformal field theory [89]. Assuming that the impurity potentials only mix states within this low-energy subspace without coupling them to higher excited levels, we first diagonalize $H_{\rm int}$ to obtain the states of delocalized quasiparticles, then diagonalize the impurity potentials in the corresponding subspace to localize quasiparticles. The advantage of this two-step method is that the translation invariance is still available in the first step to reduce the many-body Hilbert space dimension. For the $\nu=2/3$ Coulomb ground state, we find an attractive $\delta$ impurity potential
$U_{\rm imp}({\bm{w}})=-U_{0}\sum_{i=1}^{N}\delta({\bm{r}}_{i}-{\bm{w}}),$ (6)
where ${\bm{r}}_{i}$ is the position of the $i$th fermion and $U_{0}>0$, can readily pin a quasiparticle at $\bm{w}$}. The second quantization form of such an impurity potential projected to the LLL is
@@
Note that the value of $U_{0}$ cannot affect the final result of our two-step method, because only the impurity potential is involved in the second diagonalization, for which $U_{0}$ is just a prefactor.
@@ II.2 Single quasiparticle
Figure 1: (a) The density distribution for a localized single-quasiparticle excitation of the $\nu=2/3$ fermionic FQH state on a square torus with $N=25,N_{\phi}=37$. (b) The excess charge around the impurity potential. The horizontal reference line indicates $Q=e/3$. The dashed line in (b) is obtained using the first-order Haldane’s pseudopotential. Otherwise the interaction is the screened Coulomb potential.
In Fig. 1, we display the results for the largest system size, $N=25,N_{\phi}=37$, that we can reach using ED. There is a pronounced peak at the location of the impurity potential [Fig. 1(a)]. To verify that the density peak corresponds to a localized quasiparticle, we compute the excess charge
@@
where the area integral is over the disk $S_{r}$ of radius $r$ centered at the impurity potential, and $\rho_{0}({\bm{r}})$ is the density of the $\nu=2/3$ ground state without quasiparticles. In the FQH case, we simply have uniform $\rho_{0}=\nu/(2\pi\ell_{B}^{2})$. Using the fact that $\rho({\bm{r}})$ is isotropic around the impurity potential, we can simplify the excess charge as $Q(r)=2\pi e\int_{0}^{r}[\rho(r^{\prime})-\rho_{0}]r^{\prime}dr^{\prime}$, where $\rho(r)$ is the density distribution along the diagonal direction of the sample. One can see that $Q(r)$ oscillates around $e/3$ when $r\gtrsim 2\ell_{B}$ [Fig. 1(b)]. Moreover, the oscillation amplitude decays with increasing $r$. While larger system sizes are still needed for the full convergence of excess charge, our results provide compelling evidence that a $e/3$ quasiparticle is localized by the impurity potential at the sample center, which is consistent with the Abelian $\nu=2/3$ quasiparticle.
@@
$R_{\rm qp}=\sqrt{\frac{\int_{0}^{r_{\max}}|\rho(r)-\rho(r_{\max})|r^{3}dr}{\int_{0}^{r_{\max}}|\rho(r)-\rho(r_{\max})|rdr}},$ (9)
where $r_{\max}$ is the largest distance from the torus center. A numerical calculation of Eq. (9) for the system size in Fig. 1 shows $R_{\rm qp}\approx 3.9\ell_{B}$. If we replace the Coulomb interaction with the short-range first-order Haldane’s pseudopotential [91], we get $R_{\rm qp}\approx 2.6\ell_{B}$. We have also checked other system sizes with $N=19$, 2121, and $23$ particles, and get almost the same $R_{\rm qp}$ from radial density for both the screened Coulomb and pseudopotential interactions. The smaller spatial extent of the quasiparticle for the pseudopotential interaction can also be seen in the excess charge [dashed lines in Fig. 1(b)], which almost converges to $e/3$ for the same system size.
@@ II.3 Braiding of two quasiparticles
The total Berry phase can be split into two parts: one is the Aharonov-Bohm (AB) phase caused by moving a single quasiparticle in the uniform magnetic field along the same path without the other quasiparticle enclosed; and the other comes from the fractional statistics between two quasiparticles. The AB phase $p_{i}^{\rm AB}$ should be $(\pi/3)(D/\ell_{B})^{2}$, which we have numerically confirmed. Then we obtain the three fractional statistical phases as $p_{i}^{\rm anyon}=p_{i}-p_{i}^{\rm AB}$. Their averaged value $\phi_{\rm anyon}$ is displayed in Fig. 2 as a function of $D$ for $N=24,N_{\phi}=35$, which is the largest system size we can reach in the presence of two quasiparticles. With increasing $D$, the numerically obtained anyon statistical phase demonstrates the tendency of approaching the theoretical value $4\pi/3$ (equivalently, $2\pi/3$). The oscillation of $\phi_{\rm anyon}$ with $D$ reflects the oscillation of density profile of each quasiparticle. When $D$ is large enough, we expect that $\phi_{\rm anyon}$ exponentially converges to the theoretical value. However, as shown in Fig. 2, the deviation from this theoretical value still oscillates around $10^{-2}\pi$ even in the largest system size for the farthest separations between two quasiparticles. Therefore, bigger samples are required to further reduce this deviation.
The critical value of $D$ at which $\phi_{\rm anyon}$ is close enough to the theoretical value can be used as another estimation of quasiparticle size. Such a braiding-phase-based estimation is relevant to the interferometer experiment [59, 60, 61, 62, 63, 64, 65, 66, 67]. We set a threshold $10^{-2}$ for the revival in the oscillations of $|\phi_{\rm anyon}/\pi-4/3|$. Using an exponential fitting for the first two revivals, we predict the critical $D\approx 11.8\ell_{B}$, which gives the quasiparticle radius as $R_{\rm qp}\approx 5.9\ell_{B}$. Compared with the Coulomb case, the pseudopotential leads to faster convergence of the fractional statistical phase (Fig. 2), suggesting $R_{\rm qp}\approx 3.6\ell_{B}$.
It is known that quantities describing the spatial extent of a quasiparticle from different aspects may give different estimations of the quasiparticle size [90]. As the braiding phase is sensitive to the spatial overlap of two quasiparticles, it may give a larger estimation of the quasiparticle size than using the second moment of density profile. This is consistent with our numerical results: $R_{\rm qp}$ estimated by the braiding phase is $1\sim 2\ell_{B}$ larger than that from the density profile. This discrepancy is within the usual range of variation among values of $R_{\rm qp}$ obtained by distinct quantities [90].